# Optimal Control of Hybrid Systems with Submersive Resets

### 25 Jan 2024 v1

@@ Abstract
Hybrid dynamical systems are systems which posses both continuous and discrete transitions. Assuming that the discrete transitions (resets) occur a finite number of times, the optimal control problem can be solved by gluing together the optimal arcs from the underlying continuous problem via the “Hamilton jump conditions.” In most cases, it is assumed that the reset is a diffeomorphism (onto its image) and the corresponding Hamilton jump condition admits a unique solution. However, in many applications, the reset results in a drop in dimension and the corresponding Hamilton jump condition admits zero/infinitely many solutions. A geometric interpretation of this issue is explored in the case where the reset is a submersion (onto its image). Necessary conditions are presented for this case along with an accompanying numerical example.
@@
thanks: This work was funded by AFOSR Award No. MURI FA9550-32-1-0400.
@@ 1 Introduction
Here, $f\colon M\to TM$ is a smooth vector-field which encodes the continuous evolution, $\mathcal{S}\subset M$ is an embedded, codimension 1, submanifold which dictates where the discrete events occur and is called the guard, and $\Delta\colon\mathcal{S}\to M$ is the discrete event called the reset. A schematic of these dynamics where the reset fails to be injective is illustrated in Fig. 1.
@@
In the case where $\Delta$ causes a drop in dimension, i.e. $\dim(\Delta(\mathcal{S}))<\dim(\mathcal{S})$, its push-forward, $\Delta_{*}$, no longer has full rank. Resets of this type are not rare; an example of such a case is locomotion with plastic impacts, Bullo and Zefran (1998); Spong and Bullo (2005). When this happens, (4) generally admits zero or infinitely many solutions as the first condition in (4) becomes over-determined. The contribution of this work is to present a geometric solution on how to resolve this issue.
@@ 2 Optimal Control of Hybrid Systems
This work is concerned with hybrid control systems described by the tuple $\mathcal{HC}=(M,\mathcal{U},f,\mathcal{S},\Delta)$ where
@@
Let $(\bar{x},\bar{u})$ be a controlled hybrid trajectory over the interval $[t_{0},T]$ such that there exists finitely many resets and the reset times are uniformly separated. If this trajecory is optimal, then there exists a lifted hybrid curve $(\bar{x},\bar{p}):[t_{0},T]\to T^{*}M$ such that away from resets, the curve follows the Hamiltonian flow
@@ 3 Systems with Submersive Resets
A geometric interpretation of these consistency constraints is given in the following.
@@
$y(t_{0})=\bar{x}(t_{0}),\quad y(t^{*})\in\Delta^{-1}\left(\Delta(\bar{x}(t^{*}))\right).$
Intuitively, this proposition states that once the optimal trajectory is known, each individual continuous arc is optimal with known initial and terminal conditions determined by the optimal path. Call the submanifold of terminal conditions $N_{\bar{x}}\coloneqq\Delta^{-1}\left(\Delta(\bar{x}(t^{*}))\right)\subset\mathcal{S}$. The (hybrid) maximum principle applied to this restricted control problem states that the co-state has the terminal condition
@@
Assuming sufficient regularity, as a consequence of the implicit function theorem, the set of all possible solutions forms a manifold of known dimension.
@@ 4.1 Computing the Admissible Co-States
Figure 5: A plot of $\Xi_{(0,1,1)}$ along with $\tilde{\Delta}(0,-1,0.1,1,1,0)$. In this example, there exists a single intersection point.
@@
Figure 6: The optimal trajectories for 0, 1, 2, and 3 bounces.

### 19 Mar 2024 v2

@@ Abstract
Hybrid dynamical systems are systems which posses both continuous and discrete transitions. Assuming that the discrete transitions (resets) occur a finite number of times, the optimal control problem can be solved by gluing together the optimal arcs from the underlying continuous problem via the “Hamilton jump conditions.” In most cases, it is assumed that the reset is a diffeomorphism (onto its image) and the corresponding Hamilton jump condition admits a unique solution. However, in many applications, the reset results in a drop in dimension and the corresponding Hamilton jump condition admits zero/infinitely many solutions. A geometric interpretation of this issue is explored in the case where the reset is a submersion (onto its image). Optimality conditions are presented for this type of reset along with an accompanying numerical example.
@@
thanks: This work was funded by AFOSR Award No. MURI FA9550-23-1-0400.
@@ 1 Introduction
Here, $f\colon M\to TM$ is a smooth vector-field which encodes the continuous evolution, $\mathcal{S}\subset M$ is an embedded, codimension 1, submanifold which dictates where the discrete events occur and is called the guard, and $\Delta\colon\mathcal{S}\to M$ is the discrete event called the reset. A trajectory following (1) is called a hybrid arc. A schematic of these dynamics where the reset fails to be injective is illustrated in Fig. 1.
@@
In the case where $\Delta$ causes a drop in dimension, i.e. $\dim(\Delta(\mathcal{S}))<\dim(\mathcal{S})$, its push-forward, $\Delta_{*}$, no longer has full rank. Resets of this type are not rare; an example of such a case is locomotion with plastic impacts, Bullo and Zefran (1998); Spong and Bullo (2005). When this happens, (4) generally admits zero or infinitely many solutions as the first condition in (4) becomes over-determined. The contribution of this work is to present a geometric solution on how to resolve this issue. This culminates in Algorithm 3.1 which offers a way to systematically find a correct value for the co-state jumps.
@@ 2 Optimal Control of Hybrid Systems
This work is concerned with hybrid control systems of the form (2). To be explicit on the assumptions of this system, they are described by the tuple $\mathcal{HC}=(M,\mathcal{U},f,\mathcal{S},\Delta)$ where
@@
Let $(\bar{x},\bar{u})$ be a controlled hybrid trajectory over the interval $[t_{0},T]$ such that there exists finitely many resets and the reset times are uniformly separated. If this trajectory is optimal, then there exists a lifted hybrid curve $(\bar{x},\bar{p}):[t_{0},T]\to T^{*}M$ such that away from resets, the curve follows the Hamiltonian flow
@@ 3 Systems with Submersive Resets
A geometric interpretation of these consistency constraints is given in the following proposition.
@@
$y(t_{0})=\bar{x}(t_{0}),\quad y(t^{*})\in N_{\bar{x}}\coloneqq\Delta^{-1}\left(\Delta(\bar{x}(t^{*}))\right).$
Proof.
As this entire manifold, $N_{\bar{x}}$, gets mapped to the same point under $\Delta$, the location of impact on this manifold does not change the resulting trajectory. ∎
Intuitively, this proposition states that once the optimal trajectory is known, each individual continuous arc is optimal with known initial and terminal conditions determined by the optimal path. The (hybrid) maximum principle applied to this restricted control problem states that the co-state has the terminal condition
@@
Assuming sufficient regularity, the set of all possible solutions forms a manifold of known dimension.
@@
Proof.
his follows directly from the implicit function theorem. ∎
@@ 4.1 Computing the Admissible Co-States
Figure 5: A plot of $\Xi_{(0,1,1)}$ (the yellow surface) along with $\tilde{\Delta}(0,-1,0.1,1,1,0)$ (the blue curve). In this example, there exists a single intersection point.
@@
Figure 6: The optimal trajectories for 0, 1, 2, and 3 bounces (blue, red, green, and black respectively).